Please help with this problem...it's a handful. Any help is appreciated. Thanks.
Question:

Solve for y
3(y+7)/7y - (4y-7)/6 = (7y+10)/3y - 2(7y-1)/21

Found 2 solutions by stanbon, nadinadan:
Answer by stanbon(75887)

Solve for y
3(y+7)/7y - (4y-7)/6 = (7y+10)/3y - 2(7y-1)/21
Multiply thru by 42y to get rid of the denominators:
18(y+7) - 7y(4y-7) = 14(7y+10) - 4y(7y-1)
18y+ 126 -28y^2 + 49y = 98y+140 -28y^2+4y
67y+126 = 102y+140
26 = 35y
y = 26/35
=============
Cheers,
Stan H.

Answer by nadinadan(32)

So, let's find the common denominator for all fractions. 42y is the one so we multiply the first fraction by 6, the second fractions by 7y, the third by 14 and the fourth by 2y.
18(y+7) -7y(4y-7)- 14(7y+10)+4y(7y-1) / 42y Simplify
(18y+126-28y^2+49y-98y-140+28y^2-4y)/ 42y Simplify more
(70y-14)/42y -> (10y-2)/6y is the answer